OpenStudy (anonymous):

e^(-x^2) interval of convergence of expansion in power of x is (-?,?)

OpenStudy (turingtest):

what is the taylor series of e^(-x^2) ?

OpenStudy (anonymous):

it's \[\sum_{n=0}^{\infty}((-1^{n})(x ^{2n}))\div(n!)\]

OpenStudy (turingtest):

good\[\sum_{n=0}^\infty{(-1)^nx^{2n}\over n!}\] now apply the ratio test for convergence

OpenStudy (turingtest):

do you know how to do that?

OpenStudy (anonymous):

i'm looking my textbook now\

OpenStudy (turingtest):

we want to check when\[\lim_{n\to\infty}\left|{a_{n+1}\over a_n}\right|<1\]

OpenStudy (anonymous):

n! is infinite and i think x^(2n) is infinite, too

OpenStudy (anonymous):

so the answer should be (-infinite, infinite) ?

OpenStudy (anonymous):

i got -(x^2)/(n+1) what is the answer?

OpenStudy (turingtest):

yes yo have it because the limit coverges to zero the radius of covergence is infinite, and the interval of valididty is \((-\infty,\infty)\)

OpenStudy (anonymous):

thank you !